A metal cube with a mass of 55grams at temperature of 85°C is immersed in 150grams of water at temperature of 20°C. The metal and water then reach a final temperature of 23oC. What is the specific heat of the metal?
1
Expert's answer
2020-08-31T12:47:56-0400

Write the equation for thermal equilibrium. Heat the metal transferred to water is the heat water received:


"c_mm_m\Delta t_m+c_wm_w\Delta t_w=0,\\\space\\
c_m=-c_w\frac{m_w\Delta t_w}{m_m\Delta t_m},\\\space\\
c_m=-4200\frac{150(23-20)}{55(23-85)}=554\text{ J/(kg K)}."